OpenStudy (anonymous):

Can anyone help..I am trying to solve by factoring.. 3x^5/2 - 6x^3/2 = 9x^1/2

OpenStudy (anonymous):

I attempted that, but some where along the way I am getting the wrong solutions. If I factor out 3x^1/2, I get x^5 -2x^3 -3..so where do i go from there..I am not so much concerned about the answer..but how to figure it out.

OpenStudy (anonymous):

it is because factoring x^(1/2) actually means x^(5/2-1/2)

OpenStudy (anonymous):

\[x ^{a+b}=x ^{a} \times x ^{b}\]

OpenStudy (anonymous):

so then if i factor out the 3x^1/2 what should it be? 3x^2-2x^1 -3^1?

OpenStudy (anonymous):

i mean 3x^2 - 2x^1 - 3^0

OpenStudy (anonymous):

first coefficient should be 1

OpenStudy (anonymous):

1x^2 - 2x^1 - 3x^0

OpenStudy (anonymous):

why does the x stay with the 3x^0?

OpenStudy (anonymous):

x^0= 1 it was just to note that I actually did x^(0.5-0.5)

OpenStudy (anonymous):

so then i have x^2 - 2x^1 -3? which factors to (x+1(x-3) ..and the resulting factorization is 3x^(1/2) (x+1)(x-3) and the solutions are 0, -1, and 3?

OpenStudy (anonymous):

yep, confirm with wolfram if you are not sure

OpenStudy (anonymous):

the solutions in the back of the book say X=0, and x=3.... how come -1 is not a solution?

OpenStudy (anonymous):

because it is not defined in x^(0.5)

OpenStudy (anonymous):

to be more clear, the solution -1 involves the use of complex numbers

OpenStudy (anonymous):

with this kind of problems, and the ones with logarithms always go back to the original formula anc check the domain

OpenStudy (anonymous):

how would i check the domain given just the original formula

OpenStudy (anonymous):

in general things with broken exponents can't be negative 5/2 1/2, etc

OpenStudy (anonymous):

so basically just plug in the solutions? and because a sq root of a negative number is imaginary i should exclude that from the solution? it didnt give a specific domain

OpenStudy (anonymous):

yes, unles specified that roots of negative numbers are allowed, the answers must compute real

OpenStudy (anonymous):

ok, thank you so much for helping me work out this problem.. the rational exponents totally messed me up.

OpenStudy (anonymous):

np, have a nice evening or day or night,

OpenStudy (anonymous):

thankk you:) its evening...you too